# Common Trends and Long-Run Multipliers in Nonlinear Structural VARs

### v1

@@ 6 Without the common row space condition
Now suppose we were to repeat the analysis of 4 for this model: i.e. fixing the state $z_{\tau-1}=\mathfrak{z}\in\mathbb{R}^{p}$ of the model in time $\tau$, having the model impacted by a final shock $u_{\tau}=u$, and then computing the limit $\lim_{t\rightarrow\infty}z_{t}(u;\mathfrak{z})=z_{\infty}(u;\mathfrak{z})$ in the absence of any further shocks (so $u_{t}=0$ for all $t\geq\tau+1$). We may then ask: what values of $u$ would leave $z_{\infty}(u;\mathfrak{z})$ unchanged? That is, we would like to determine the set
@@
Let us suppose (unlike what is maintained in 4 above), that $\beta^{\mathsf{T}}z_{\tau-1}=\beta^{\mathsf{T}}\mathfrak{z}=0$, so that $z_{\tau-1}\in\mathscr{M}_{0}$, and the model is in equilibrium prior to the incidence of $u_{\tau}$. Locally to $\mathscr{M}_{0}$, we would expect shocks in the direction of $\operatorname{sp}\tilde{\alpha}$ to have no permanent effects; whereas further from $\mathscr{M}_{0}$, shocks in directions lying progressively closer to $\operatorname{sp}\alpha$ should have no permanent effects.
@@
and $\Lambda(\xi)=2\lvert\Phi(\xi)-0.5\rvert$. (Without loss of generality, we take $\mathfrak{z}=0$.) For each $\lVert u\rVert$ between 00 and $10$, the figure reports the (unique) direction $\delta_{\lVert u\rVert}$ such that $u_{\tau}=\lVert u\rVert\delta_{\lVert u\rVert}$ has no permanent effect on $z_{t}$; for the sake of comparability with $\tilde{\alpha}$ and $\alpha$, both of which have unit first element, this is reported in terms of the ratio $\delta_{\lVert u\rVert,2}/\delta_{\lVert u\rVert,1}$. We see here that $\delta_{\lVert u\rVert,2}/\delta_{\lVert u\rVert,1}$ takes the value $0.5=\tilde{\alpha}_{2}$ when $\lVert u\rVert=0$, and tends towards $0.25=\alpha_{1}$ as $\lVert u\rVert$ grows (equalling $0.26$ when $\lVert u\rVert=20$). As a consequence
@@
and as such there is no direction along which the reduced form shocks $u_{t}$ will only have transient effects. There is thus no possibility of discriminating between the underlying structural shocks $\varepsilon_{t}$ according to whether not they have permanent effects: each must, depending on the magnitude of the shock.
7 Piecewise affine VARs
@@
$\mathbf{1}^{(\ell)}(z)\coloneqq\mathbf{1}\{z\in\mathscr{Z}_{i}^{(\ell)}\},$
@@ 7.2 Membership of ${\cal M}_{r}$
Since it may be verified that a linear VAR satisfying (i)–3 is itself a piecewise linear model satisfying (i)–3 above, 2.1 may be construed as a special case of the preceding (though for expository reasons, a separate proof of that result is given in D).
@@ 7.3 Smooth transitions
where $\Lambda^{(\ell)}(z)\in[0,1]$ and $\sum_{\ell=1}^{L}\Lambda^{(\ell)}(z)=1$ for all $z\in\mathbb{R}^{p}$, so that $f_{i}^{\mathrm{ST}}(z)$ is a always a smooth, convex combination of the affine functions $\{z\mapsto\bar{\phi}_{i}^{(\ell)}+\Phi_{i}^{(\ell)}z\}_{\ell=1}^{L}$. While models of this kind may be accommodated within our framework (under certain regularity conditions), the fact that the gradient of $f_{i}^{\mathrm{ST}}$ is not a convex combination of those underlying affine regimes makes it difficult to reduce (iii) to a bound on the JSR of a finite collection of matrices, in the manner of 7.1. As an alternative specification that allows for smooth transitions between regimes, while also keeping (iii) tractable, consider

### v2

@@ 5 The common row space condition
Now suppose we were to repeat the analysis of 4.1 for this model: i.e. fixing the state $z_{\tau-1}=\mathfrak{z}\in\mathbb{R}^{p}$ of the model in time $\tau$, having the model impacted by a final shock $u_{\tau}=u$, and then computing $\lim_{t\rightarrow\infty}z_{t}(u;\mathfrak{z})=z_{\infty}(u;\mathfrak{z})$ in the absence of any further shocks (so $u_{t}=0$ for all $t\geq\tau+1$). We may then ask: what values of $u$ would leave $z_{\infty}(u;\mathfrak{z})$ unchanged? That is, we would like to determine the set
@@
Let us further suppose that $\beta^{\mathsf{T}}z_{\tau-1}=\beta^{\mathsf{T}}\mathfrak{z}=0$, so that $z_{\tau-1}\in\mathscr{M}_{0}$, and the model is in equilibrium prior to the incidence of $u_{\tau}$. Locally to $\mathscr{M}_{0}$, we would expect shocks in the direction of $\operatorname{sp}\tilde{\alpha}$ to have no permanent effects; whereas further from $\mathscr{M}_{0}$, shocks in directions lying progressively closer to $\operatorname{sp}\alpha$ should have no permanent effects.
@@
and $\Lambda(\xi)=2\lvert\Phi(\xi)-0.5\rvert$, where $\Phi$ denotes the Gaussian cdf. (Without loss of generality, we take $\mathfrak{z}=0$.) For values of $\lVert u\rVert$ between 00 and $10$, the figure reports the (unique) direction $\delta_{\lVert u\rVert}$ such that $u_{\tau}=\lVert u\rVert\delta_{\lVert u\rVert}$ has no permanent effect on $z_{t}$; for the sake of comparability with $\tilde{\alpha}$ and $\alpha$, both of which have unit first element, this is reported in terms of the ratio $\delta_{\lVert u\rVert,2}/\delta_{\lVert u\rVert,1}$. We see here that $\delta_{\lVert u\rVert,2}/\delta_{\lVert u\rVert,1}$ takes the value $0.5=\tilde{\alpha}_{2}$ when $\lVert u\rVert=0$, and tends towards $0.25=\alpha_{1}$ as $\lVert u\rVert$ grows (equalling $0.26$ when $\lVert u\rVert=20$). As a consequence
@@
and as such there is no direction along which the shocks $u_{t}$ will only have transient effects. There is thus no possibility of discriminating between the underlying structural shocks $\varepsilon_{t}$ according to whether not they have permanent effects: each must have some such effect, to an extent that varies with the magnitude of the shock.
6 Application to piecewise affine VARs
@@
$\mathbf{1}^{(\ell)}(z)\coloneqq\mathbf{1}\{z\in\mathscr{Z}^{(\ell)}\},$
@@ 6.2 Membership of ${\cal M}_{r}$
Since it may be verified that a linear VAR satisfying (i)–3 is itself a piecewise linear model satisfying (i)–3 above, 2.1 may be construed as a special case of the preceding (though for expository reasons, a separate proof of that result is given in E).
@@ 6.3 Smooth transitions
where $\Lambda^{(\ell)}(z)\in[0,1]$ and $\sum_{\ell=1}^{L}\Lambda^{(\ell)}(z)=1$ for all $z\in\mathbb{R}^{p}$, so that $f_{i}^{\mathrm{ST}}(z)$ is a always a smooth, convex combination of the affine functions $\{z\mapsto\bar{\phi}_{i}^{(\ell)}+\Phi_{i}^{(\ell)}z\}_{\ell=1}^{L}$. While models of this kind may be accommodated within our framework (under certain regularity conditions), the fact that the gradient of $f_{i}^{\mathrm{ST}}$ is not a convex combination of those underlying affine regimes makes it difficult to reduce (iii) to a bound on the JSR of a finite collection of matrices, in the manner of 6.1. As an alternative specification that allows for smooth transitions between regimes, while also keeping (iii) tractable, consider